OpenStudy (anonymous):

Write an equation for the line in slope-intercept form, given that the line passes through the point (2,3) and (3,4).

OpenStudy (anonymous):

?

hero (hero):

Have you found the slope yet?

OpenStudy (anonymous):

No, that's why I asked this question. I don't really understand slope.

OpenStudy (anonymous):

me neither was wanting to see what the answer was2 sorry

hero (hero):

Hmmm

hero (hero):

Well, first of all the points given have the coordinates (x,y)

hero (hero):

There is a formula for slope based on the coordinates of the given points that is used to find the slope: \[\frac {y_2 - y_1}{x_2 - x_1}\]

hero (hero):

If you take the coordinates of each point and label them like this: \[x_1 =2 \] \[ y_1= 3\] \[ x_2 = 3\] \[y_2 = 4\] You can substitute those numerical values into the formula and find the slope

hero (hero):

\[\frac{4-3}{3-2} = 1\] So the slope is just 1

hero (hero):

I don't have time to finish helping you with this, sorry.